OpenStudy (anonymous):

If the equation of a circle is (x + 5)^2 + (y - 7)^2 = 36, its radius is: A. 6 B. 16 C. 36

OpenStudy (welshfella):

compare your equation with the general form (x - a)^2 + (y - b)^2 = r^2 where the center has coordinates (a,b) and r = radius

OpenStudy (welshfella):

line them up and you'll see what the radius is

OpenStudy (anonymous):

the radius is 6 then correct?

OpenStudy (welshfella):

yes 6^2 = r^2 = 36

OpenStudy (anonymous):

thank you :)

OpenStudy (welshfella):

yw